# Physical insights from the aspect ratio dependence of turbulence in negative triangularity plasmas

### v1

@@ Abstract
In this work, we study the impact of aspect ratio $A=R_{0}/r$ (the ratio of major radius $R_{0}$ to minor radius r) on the confinement benefits of Negative Triangularity (NT) plasma shaping. We use high-fidelity flux tube gyrokinetic GENE simulations and consider several different scenarios: four of them inspired by TCV experimental data, a scenario inspired by DIII-D experimental data and a scenario expected in the new SMART spherical tokamak. The present study reveals a surprising and non-trivial dependence. NT improves confinement at any value of A for ITG turbulence, while for TEM turbulence confinement is improved only in the case of large and conventional aspect ratios. Additionally, through a detailed study of a large aspect ratio case with pure ITG drive, we develop an intuitive physical picture that explains the beneficial effect of NT at large and conventional aspect ratios. This picture does not hold in TEM-dominated regimes, where a complex synergistic effect of many factors is found. Finally, we performed the first linear gyrokinetic simulations of SMART, finding that both NT and PT scenarios are dominated by micro-tearing-mode (MTM) turbulence and that NT is more susceptible to MTMs at tight aspect ratio. However, we found that a regime where ITGs dominate in SMART can be found, and in this regime NT is more linearly stable.
@@ I Introduction
The discovery of the High confinement mode (H-mode) in ASDEX [1] opened the gates to a new era of research in the field of magnetically confined fusion plasmas. After exceeding a certain threshold of heating power, a tokamak plasma spontaneously creates a transport barrier at the edge, thereby enabling the formation of a region of steep pressure gradient. This is the so-called pedestal which greatly increases the temperature and density in the core of the plasma. The performance of H-mode plasmas is substantially better than standard Low confinement mode (L-mode) plasmas, which do not display a noticeable pedestal. However, the promise of H-mode is tempered by the appearance of powerful instabilities called Edge Localized Modes (ELMs), which are quasi-periodic relaxations of the pedestal that release large amounts of particles and energy into the plasma-facing components of the tokamak. In addition, the power released by an ELM is directly proportional to the power injected in the plasma, making ELMs extremely dangerous for the integrity of a power plant. A single ELM could seriously damage the first wall of the machine and compromise the steady production of energy.
For these reasons, the European roadmap towards DEMO (DEMOnstration power plant) identified the ELMs as one of the most important issues to be solved [2] and recommends searching for methods to reduce the size of ELMs or for scenarios where they can be avoided without compromising confinement. A solution may come from plasma shaping.
Decades of plasma physics research have proven that giving a tokamak plasma a non-circular poloidal cross-sectional shape can greatly modify the stability of the plasma and its confinement of particles and energy. Figure 1 shows how the poloidal cross-sectional shape of a plasma can be modified by changing the two most important shaping factors: elongation
@@
The configuration that has become the standard scenario in the majority of modern tokamaks is the elongated dee-shaped one, also called Positive Triangularity (PT) because the value of $\delta$ is larger than zero. The term dee comes from the resemblance of the cross-section to the letter ”D”, when shown to the right of the tokamak axis of symmetry (as is typical). PT plasmas became popular because they have good MHD stability and can enter H-mode with less heating power. Of course, the downside is the aforementioned ELMs, the seriousness of which has been widely appreciated only more recently. An alternative that has gained interest is Negative Triangularity (NT), where the plasma is shaped such that $\delta$ is less than zero. A large number of experiments on several medium-size tokamaks (on TCV for several decades [3, 4, 5], followed more recently by DIII-D [6, 7] and AUG [8]) as well as many numerical studies [9, 10, 11], find that NT reduces turbulent transport, and can inhibit the transition to H-mode. These features allow a NT plasma to have an L-mode-like edge pressure profile and avoid ELMs, but with sufficiently increased pressure gradient at the edge and in the outer core to recover H-mode-like values of density and temperature in the core.
Figure 1: The effect of (a) elongation $\kappa$ with $\delta=0$ and (b) triangularity $\delta$ with $\kappa=1$ on the poloidal cross-section.
@@
Despite the great number of studies on the topic, the physical processes underlying the beneficial effect of NT on turbulent transport are not fully understood yet. Moreover, as observed by Merlo and Jenko in [10], it is key to study how other plasma parameters influence the impact of NT on confinement so we can use this knowledge to optimize. In this work, we will study the dependence of transport on the aspect ratio $A=R_{0}/r$ (the ratio of major radius $R_{0}$ to minor radius $r$) in NT plasmas. The motivation for such a study is twofold. First, studying how a NT plasma behaves at low A gives insight into the feasibility of a NT Spherical Tokamak (ST), i.e. a tokamak characterized by a small aspect ratio. Second, studying the large aspect ratio limit greatly simplifies the physics and therefore the analysis. Moreover, as we will see, the conclusions will turn out to be valid also at the aspect ratios typical for a standard tokamak.
As the aspect ratio is not a parameter that can be easily changed and scanned in real experiments, we make use of high-fidelity flux tube gyrokinetic GENE [12] simulations, which allow us to explore the parameter space and capture the physical processes controlling turbulence. To carry out a comprehensive analysis, we will consider scenarios characterized by different turbulent regimes, i.e. regimes with different dominant turbulent drive. Moreover, to maintain realism, we will use discharges produced in TCV and DIII-D. In total, we will consider 5 different cases: a realistic Trapped Electron Mode (TEM) dominated scenario (rTEM), a realistic Ion Temperature Gradient (ITG) dominated scenario (rITG), a pure density gradient-driven TEM scenario (pTEM), and two pure ion temperature gradient-driven scenarios (pITG-1 and pITG-2). All the simulations are performed using the Miller local equilibrium model [13] to construct local analytic magnetic equilibria around a specific flux surface, which allows us to then change the aspect ratio as well as the sign of triangularity. Moreover, to isolate the effect of geometry, the turbulence drives, namely the logarithmic gradients of density and temperature, are kept fixed.
@@
In this paper, we will see that a non-trivial picture will emerge from the aspect ratio scan, showing that NT has better confinement than PT at large and standard aspect ratio values, regardless of the nature of the turbulent regime. However, at tight A NT becomes detrimental in the TEM-dominated regime, while staying beneficial in the ITG-dominated regime. This complex dependency on the aspect ratio triggered a deeper study to understand the physical processes underlying the stabilization/destabilization of certain types of turbulence by a NT geometry. While an explanation of the behaviour at small A is still missing, we will present a simple but compelling physical picture of why NT weakens ITG instability at large and standard aspect ratios. We refer to this understanding as simple because it relies only on two quantities, i.e. magnetic drifts and Finite Larmor Radius (FLR) effects, which in turn are easy to calculate from the shape of the plasma. Motivated by similarities between our model and the work done in [11] for TEM turbulence, we also tried to apply our model to a pTEM case, finding that FLR effects and magnetic drifts can only explain part of the physics for this kind of scenario.
@@
Finally, to achieve a better understanding of the small A regime and also use more realistic equilibria, we will consider the ST SMART [14, 15, 16, 17]. This is a ST that only very recently started operation at the University of Seville, characterized by strong geometric flexibility, which should enable the creation of single-null and double-null NT and PT configurations. The input parameters for GENE simulations are taken from predictive simulations for SMART made with TRANSP [18]. We will perform only local linear simulations. They will reveal which kind of turbulence dominates and how it responds to a change in various parameters. We will see that, because of the high $\beta$ typical of STs, large electron temperature gradient and large magnetic shear, turbulence is dominated by Micro-Tearing Modes (MTM). Moreover, we will find that NT is more sensitive to MTMs, with growth rates larger than those computed for PT. However, we will see that it should be possible to operate SMART in an ITG-dominated regime, where it is possible to exploit the beneficial effect of NT (consistently with previous simulations). We were not able to perform reliable NL simulations because of the challenges commonly encountered when simulating scenarios mainly driven by MTMs (e.g. lack of converged heat fluxes when using standard resolutions). This is left as future work.
@@
The remainder of this paper is organized as follows. In section II, we describe briefly how GENE works and how it has been used to carry out the present study. In section III, we present how the aspect ratio and turbulent regime can modify the beneficial effect of NT on confinement. In section IV, we motivate and propose a simple physical picture of why NT weakens ITG turbulence in large and standard aspect ratio tokamaks. In section V, we apply the model described in the previous section to a pTEM scenario. In section VI, we present the first linear simulations performed for the SMART tokamak, providing predictions for the turbulent regime and the performance of the NT option with respect to PT. Finally, section VII draws some conclusions and provides a final discussion.
@@ II Numerical model and methodology
This work consists exclusively of numerical simulations performed with the flux tube gradient-driven version of the GENE code [12]. GENE is a physically comprehensive Eulerian gyrokinetic code that solves the Vlasov-Maxwell equations discretized on a 5-dimensional (5D) grid. This leads to a large set of coupled ordinary differential equations that can be solved as an initial value problem or as an eigenvalue problem. The code can be run in a linear mode if the nonlinear terms of the GK equation are neglected, or in a nonlinear mode when the nonlinear terms are retained. The former is considerably faster and is especially useful to quickly gain an understanding of the nature of turbulence. The latter is computationally more expensive but needed to physically model micro-turbulence. The real space is parametrized by a 3D system of non-orthogonal field-aligned coordinates $(x,y,z)$, which correspond respectively to the radial, binormal, and parallel (to the magnetic field B) coordinates. They are defined by
@@
where $\psi$ is the poloidal magnetic flux, $\varphi$ the toroidal angle, $\theta$ the straight field-line poloidal angle, $q$ is the safety factor and $C_{y}$ is a constant. The remaining two coordinates of the 5D discretized grid, $v_{\parallel}$ and $\mu$, describe velocity space and correspond to the parallel velocity and the magnetic moment (defined as $\mu=(m_{\sigma}v_{\perp}^{2})/(2B)$, where $m_{\sigma}$ is the mass of the spiecies $\sigma$ and $v_{\perp}$ the perpendicular velocity). In a flux tube, the background values of density, temperature, and flow (as well as their gradients) are evaluated at the flux surface of interest and kept constant across the simulation domain. This choice is justified if the radial scale of turbulence is much smaller than the radial scale of variation of equilibrium quantities (machine scale), which is fairly well satisfied for TCV and better satisfied in larger devices like DIII-D or reactor-scale devices. The real space simulation domain corresponds to a small rectangle extended in $x$ and $y$, that follows the magnetic field lines along $z$. Periodic boundary conditions are used for the perpendicular coordinates $x$ and $y$, while pseudo-periodic boundary conditions are applied to $z$. For more details, the interested reader is referred to [19].
GENE can use different methods or codes to reconstruct the magnetic equilibrium at a specific flux surface. These can be divided into numerical reconstruction codes (CHEASE [20], gist, and tracer-efit) and analytical models ($s-\alpha$ and Miller). For all the simulations presented in this work, we used the local equilibrium Miller model, because, as an analytical model, it allows us to easily change the geometry by changing a few parameters. As shown in [21], a Miller equilibrium is completely defined by a total of 14 scalar quantities. As shown in equation 4, the parametrization of the flux surface shape requires 6 parameters: the geometric axis $R_{0}$, the elevation $Z_{0}$, the aspect ratio $A=R_{0}/r$, the elongation $\kappa$, the triangularity $\delta$ and the squareness $\zeta$.
@@
To perform our simulations, we used Miller to model relevant flux surfaces in actual experimental magnetic equilibria produced on TCV and DIII-D. We flipped the triangularity (from NT to PT or vice versa) and changed the aspect ratio varying the minor radius. During these operations, the values of density and temperature, and their logarithmic gradients, were kept fixed. This procedure is commonly used in GK simulations and enables one to isolate the effect of geometry on transport. The numerical convergence of all simulations has been tested by resolution studies. More details will be given in the next sections, but, depending on the simulation, we used $L_{x}\sim[110-180]\rho_{i}$ and $L_{y}\sim[125-250]\rho_{i}$ for the radial and binormal widths of the simulation box with $[128-256]$ $k_{x}$ and $[32-64]$ $k_{y}$ modes. Depending on the aspect ratio, we used between 32 to 48 grid points in $z$. The size and resolution of velocity space have been kept fixed: $L_{v}=3v_{th,\sigma}$, $L_{\mu}=9\mu B_{0}/T_{\sigma}$, $n_{v}=42$ and $n_{\mu}=12$, where $v_{th,\sigma}=\sqrt{2T_{\sigma}/m_{\sigma}}$ is the thermal velocity, $B_{0}$ the toroidal magnetic field at $R_{0}$ on the flux surface of interest, $m_{\sigma}$ the reference mass and $T_{\sigma}$ the reference temperature.
@@ III Aspect ratio scan
As mentioned in section I, to carry out the most realistic analysis possible, we based our parameters on an Electron-Cyclotron-Resonance (ECR) heated TCV NT discharge (#68145), a Neutral Beam Injection (NBI) heated TCV NT discharge (#69682), and a high power NBI-heated DIII-D PT discharge (#176492). We reconstructed the magnetic equilibria with Miller, choosing a radius of analysis (defined here as the normalized square root of the toroidal flux) $\rho_{tor}=0.8$ for TCV and $\rho_{tor}=0.75$ for DIII-D. Then we flipped the sign of $\delta$ and $s_{\delta}$ and we artificially changed the aspect ratio keeping fixed the major radius and changing the minor radius. At the chosen radial locations, the logarithmic gradients were computed from the experimental data and then were kept fixed while the geometry was changed.
@@
We chose discharges #68145 and #176492 to have cases dominated by TEM and ITG turbulence respectively. Originally, #69682 was chosen to provide an ITG-dominated case, but linear GK simulations revealed that it was TEM-dominated. Additionally, we modified the logarithmic gradients of #68145 and #69682 to create idealized cases, namely two pure ion temperature gradient-driven scenarios (pITG-1 and pITG-2) and a pure density gradient-driven TEM one (pTEM). Some of the most important input parameters used in the simulations are reported in table 1, while the shapes of the three original scenarios with their flipped counterparts are shown in figure 2. For the reaistic TCV and DIII-D simulations collisions, electromagnetic effects and carbon as the main impurity were retained. The pITG and pTEM simulations neglected impurities, but collisions and electromagnetic effects were kept. All simulations used a kinetic treatement for the electrons.
@@
(a) pITG-1
(b) pITG-2
(c) rITG
(d) rTEM
Heat flux spectra are plotted in figure 4 for different regimes using two values of aspect ratio: conventional and spherical. We are not showing the spectra for the large aspect ratio cases because they are similar to conventional A. We show only the most important component, i.e. electron or ion channel, depending on the considered regime. From figures 4(a), 4(b) and 4(c), which correspond to ITG regimes, we can see that at conventional A the reduction of heat flux is equally distributed among all the considered $k_{y}$ modes. Instead, when we move to a smaller A, the reduction is more pronounced for modes with $k_{y}\rho_{s}$ bigger than 0.2. If we now look at figure 4(d), which corresponds to the realistic TEM case, we can see that the detrimental effect of NT comes from small $k_{y}$ modes at tight A. These are the modes that are most strongly destabilized when one decreases the aspect ratio for TEM turbulence.

### v2

@@ Abstract
In this work, we study the impact of aspect ratio $A=R_{0}/r$ (the ratio of major radius $R_{0}$ to minor radius r) on the confinement benefits of Negative Triangularity (NT) plasma shaping. We use high-fidelity flux tube gyrokinetic GENE simulations and consider several different scenarios: four of them inspired by TCV experimental data, a scenario inspired by DIII-D experimental data and a scenario expected in the new SMART spherical tokamak. The present study reveals a distinct and non-trivial dependence. NT improves confinement at any value of $A$ for ITG turbulence, while for TEM turbulence confinement is improved only in the case of large and conventional aspect ratios. Additionally, through a detailed study of a large aspect ratio case with pure ITG drive, we develop an intuitive physical picture that explains the beneficial effect of NT at large and conventional aspect ratios. This picture does not hold in TEM-dominated regimes, where a complex synergistic effect of many factors is found. Finally, we performed the first linear gyrokinetic simulations of SMART, finding that both NT and PT scenarios are dominated by micro-tearing-mode (MTM) turbulence and that NT is more susceptible to MTMs at tight aspect ratio. However, a regime where ITG dominates can be found in SMART, and in this regime NT is more linearly stable.
@@ I Introduction
The discovery of the High confinement mode (H-mode) in ASDEX Team (1989) opened the gates to a new era of research in the field of magnetically confined fusion plasmas. After exceeding a certain threshold of heating power, a tokamak plasma spontaneously creates a transport barrier at the edge, thereby enabling the formation of a region of steep pressure gradient. This is the so-called pedestal which greatly increases the temperature and density in the core of the plasma. The performance of H-mode plasmas is substantially better than standard Low confinement mode (L-mode) plasmas, which do not display a noticeable pedestal. However, the promise of H-mode is tempered by the appearance of powerful instabilities called Edge Localized Modes (ELMs) Leonard (2014), which are quasi-periodic relaxations of the pedestal that release large amounts of particles and energy into the plasma-facing components of the tokamak. In addition, the power released by an ELM is directly proportional to the power injected in the plasma, making ELMs extremely dangerous for the integrity of a power plant. A single ELM could seriously damage the first wall of the machine and compromise the steady production of energy.
For these reasons, the European roadmap towards DEMO (DEMOnstration power plant) identified the ELMs as one of the most important issues to be solved EUROfusion and recommends searching for methods to reduce the size of ELMs or for scenarios where they can be avoided without compromising confinement. A solution may come from plasma shaping.
Decades of plasma physics research have proven that giving a tokamak plasma a non-circular poloidal cross-sectional shape can greatly modify the stability of the plasma and its confinement of particles and energy. The poloidal cross-sectional shape of a plasma can be modified by changing the two most important shaping factors: elongation
@@
The configuration that has become the standard scenario in the majority of modern tokamaks is the elongated dee-shaped one, also called Positive Triangularity (PT) because the value of $\delta$ is larger than zero. The term dee comes from the resemblance of the cross-section to the letter ”D”, when shown to the right of the tokamak axis of symmetry (as is typical). PT plasmas became popular because they have good MHD stability and can enter H-mode with less heating power. Of course, the downside is the aforementioned ELMs, the seriousness of which has been widely appreciated only more recently. An alternative that has gained interest is Negative Triangularity (NT), where the plasma is shaped such that $\delta$ is less than zero. A large number of experiments on several medium-size tokamaks (on TCV for several decades Pochelon et al. (1998); Fontana et al. (2017); Coda et al. (2021), followed more recently by DIII-D Austin et al. (2018); Marinoni et al. (2019) and AUG Happel et al. (2022)) as well as many numerical studies Merlo et al. (2021); Merlo and Jenko (2023); Marinoni et al. (2009), find that NT reduces turbulent transport, and can inhibit the transition to H-mode. These features allow a NT plasma to have an L-mode-like edge pressure profile and avoid ELMs, but with sufficiently increased pressure gradient at the edge and in the outer core to recover H-mode-like values of density and temperature in the core.
@@
Despite the great number of studies on the topic, the physical processes underlying the beneficial effect of NT on turbulent transport are not fully understood yet. Moreover, as observed by Merlo and Jenko in Merlo and Jenko (2023), it is key to study how other plasma parameters influence the impact of NT on confinement so we can use this knowledge to optimize. In this work, we will study the dependence of transport on the aspect ratio $A=R_{0}/r$ (the ratio of major radius $R_{0}$ to minor radius $r$) in NT plasmas. The motivation for such a study is twofold. First, studying how a NT plasma behaves at low $A$ gives insight into the feasibility of a NT Spherical Tokamak (ST), i.e. a tokamak characterized by a small aspect ratio. Second, studying the large aspect ratio limit greatly simplifies the physics and therefore the analysis. Moreover, as we will see, the conclusions will turn out to be valid also at the aspect ratios typical for a standard tokamak. As the aspect ratio is not a parameter that can be easily changed and scanned in real experiments, we make use of high-fidelity flux tube gyrokinetic GENE Jenko et al. (2000) simulations, which allow us to explore the parameter space and capture the physical processes controlling turbulence.
@@
In this paper, we will see that a non-trivial picture will emerge from the aspect ratio scan, showing that NT has better confinement than PT at large and standard aspect ratio values, regardless of the nature of the turbulent regime. However, at tight $A$ NT becomes detrimental in Trapped Electron Mode (TEM) dominated regimes, while staying beneficial in Ion Temperature Gradient (ITG) mode dominated regimes. This complex dependency on the aspect ratio triggered a deeper study to understand the physical processes underlying the stabilization/destabilization of certain types of turbulence by a NT geometry. While an explanation of the behaviour at small $A$ is still missing, we will present a simple but compelling physical picture of why NT weakens ITG instability at large and standard aspect ratios. We refer to this understanding as simple because it relies only on two quantities, i.e. magnetic drifts and Finite Larmor Radius (FLR) effects, which in turn are easy to calculate from the shape of the plasma. Motivated by similarities between our model and the work done in Marinoni et al. (2009) for TEM turbulence, we also tried to apply our model to a scenario with pure TEM drive, finding that FLR effects and magnetic drifts can only explain part of the physics for this kind of scenario.
@@
Finally, to achieve a better understanding of the small $A$ regime and also use more realistic equilibria, we will consider the ST SMART Doyle et al. (2021a, b); Mancini et al. (2023); Agredano-Torres et al. (2021). This is a ST that only very recently started operation at the University of Seville, characterized by strong geometric flexibility, which should enable the creation of single-null and double-null NT and PT configurations. The input parameters for GENE simulations are taken from predictive simulations for SMART made with TRANSP Cruz-Zabala et al. (2024). We will perform only local linear simulations. They will reveal which kind of turbulence dominates and how it responds to a change in various parameters. We will see that, because of the high $\beta$ typical of STs, large electron temperature gradient and large magnetic shear, turbulence is dominated by Micro-Tearing Modes (MTM). Moreover, we will find that NT is more sensitive to MTMs, with growth rates larger than those computed for PT. However, we will see that it should be possible to operate SMART in an ITG-dominated regime, where it is possible to exploit the beneficial effect of NT (consistently with previous simulations). We were not able to perform reliable NL simulations because of the challenges commonly encountered when simulating scenarios mainly driven by MTMs (e.g. lack of converged heat fluxes when using standard resolutions).
@@
The remainder of this paper is organized as follows. In section II, we describe briefly how GENE works and how it has been used to carry out the present study. In section III, we present how the aspect ratio and turbulent regime can modify the beneficial effect of NT on confinement. In section IV, we present the first linear simulations performed for the SMART tokamak, providing predictions for the turbulent regime and the performance of the NT option with respect to PT. In section V, we motivate and propose a simple physical picture of why NT weakens ITG turbulence in large and standard aspect ratio tokamaks. In section VI, we apply the model described in the previous section to a scenario with pure TEM drive. Finally, section VII draws some conclusions and provides a final discussion.
@@ II Numerical model and methodology
This work consists exclusively of numerical simulations performed with the flux tube gradient-driven version of the GENE code Jenko et al. (2000). GENE is a physically comprehensive Eulerian gyrokinetic code that solves the Vlasov-Maxwell equations discretized on a 5-dimensional (5D) grid. This leads to a large set of coupled ordinary differential equations that can be solved as an initial value problem or as an eigenvalue problem. The code can be run in a linear mode if the nonlinear terms of the GK equation are neglected, or in a nonlinear mode when the nonlinear terms are retained. The former is considerably faster and is especially useful to quickly gain an understanding of the nature of turbulence. The latter is computationally more expensive but needed to physically model micro-turbulence. The real space is parametrized by a 3D system of non-orthogonal field-aligned coordinates $(x,y,z)$, which correspond respectively to the radial, binormal, and parallel (to the magnetic field B) coordinates. They are defined by
@@
where $\psi$ is the poloidal magnetic flux, $\varphi$ the toroidal angle, $\theta$ the straight field-line poloidal angle, $q$ is the safety factor and $C_{y}$ is a constant. The remaining two coordinates of the 5D discretized grid, $v_{\parallel}$ and $\mu$, describe velocity space and correspond to the parallel velocity and the magnetic moment (defined as $\mu=(m_{\sigma}v_{\perp}^{2})/(2B)$, where $m_{\sigma}$ is the mass of the spiecies $\sigma$ and $v_{\perp}$ the perpendicular velocity). In a flux tube, the background values of density, temperature, and flow (as well as their gradients) are evaluated at the flux surface of interest and kept constant across the simulation domain. This choice is justified if the radial scale of turbulence is much smaller than the radial scale of variation of equilibrium quantities (machine scale), which is fairly well satisfied for TCV and better satisfied in larger devices like DIII-D or reactor-scale devices. The real space simulation domain corresponds to a small rectangle extended in $x$ and $y$, that follows the magnetic field lines along $z$. Periodic boundary conditions are used for the perpendicular coordinates $x$ and $y$, while pseudo-periodic boundary conditions are applied to $z$. For more details, the interested reader is referred to Beer et al. (1995).
To carry out a comprehensive analysis, we will consider scenarios characterized by different turbulent regimes, i.e. regimes with different dominant turbulent drive. Moreover, to maintain realism, we will use discharges produced in TCV and DIII-D, and scenarios predicted for SMART. In total, we will consider six different cases: a realistic TEM dominated scenario (rTEM), a realistic ITG dominated scenario (rITG), a pure density gradient-driven TEM scenario (pTEM), two pure ion temperature gradient-driven scenarios (pITG-1 and pITG-2), and an MTM dominated scenario (the one relative to SMART).
GENE can use different methods or codes to reconstruct the magnetic equilibrium at a specific flux surface. These can be divided into numerical reconstruction codes (CHEASE Lütjens et al. (1996), gist, and tracer-efit) and analytical models ($s-\alpha$ and Miller). For all the simulations presented in this work, we used the local equilibrium Miller model, because, as an analytical model, it allows us to easily change the geometry by changing a few parameters. As shown in Candy (2009), a Miller equilibrium is completely defined by a total of 14 scalar quantities. As shown in equation 4, the parametrization of the flux surface shape requires 6 parameters: the geometric axis $R_{0}$, the elevation $Z_{0}$, the aspect ratio $A=R_{0}/r$, the elongation $\kappa$, the triangularity $\delta$ and the squareness $\zeta$.
@@
To perform our simulations, we used Miller to model relevant flux surfaces in actual experimental magnetic equilibria. We flipped the triangularity (from NT to PT or vice versa) and changed the aspect ratio varying the minor radius. During these operations, the values of density and temperature, and their logarithmic gradients, were kept fixed. This procedure is commonly used in GK simulations and enables one to isolate the effect of geometry on transport. The numerical convergence of all simulations has been tested by resolution studies. More details will be given in the next sections, but, depending on the simulation, we used $L_{x}\sim[110-180]\rho_{i}$ and $L_{y}\sim[125-250]\rho_{i}$ for the radial and binormal widths of the simulation box with $[128-256]$ $k_{x}$ and $[32-64]$ $k_{y}$ modes. Depending on the aspect ratio, we used between 32 to 48 grid points in $z$. The size and resolution of velocity space have been kept fixed: $L_{v}=3v_{th,\sigma}$, $L_{\mu}=9\mu B_{0}/T_{\sigma}$, $n_{v}=42$ and $n_{\mu}=12$, where $v_{th,\sigma}=\sqrt{2T_{\sigma}/m_{\sigma}}$ is the thermal velocity, $B_{0}$ the toroidal magnetic field at $R_{0}$ on the flux surface of interest, $m_{\sigma}$ the reference mass and $T_{\sigma}$ the reference temperature.
@@ III Aspect ratio scan
To carry out the most realistic analysis possible, we based our parameters on an Electron-Cyclotron-Resonance (ECR) heated TCV NT discharge (#68145), a Neutral Beam Injection (NBI) heated TCV NT discharge (#69682), and a high power NBI-heated DIII-D PT discharge (#176492). We reconstructed the magnetic equilibria with Miller, choosing a radius of analysis (defined here as the normalized square root of the toroidal flux) $\rho_{tor}=0.8$ for TCV and $\rho_{tor}=0.75$ for DIII-D. Then we flipped the sign of $\delta$ and $s_{\delta}$ and we artificially changed the aspect ratio keeping fixed the major radius and changing the minor radius. At the chosen radial locations, the logarithmic gradients were computed from the experimental data and then were kept fixed while the geometry was changed.
@@
We chose discharges #68145 and #176492 to have cases dominated by TEM and ITG turbulence respectively. Originally, #69682 was chosen to provide an ITG-dominated case, but linear GK simulations revealed that it was TEM-dominated. Additionally, we modified the logarithmic gradients of #68145 and #69682 to create idealized cases, namely two pure ion temperature gradient-driven scenarios (pITG-1 and pITG-2) and a pure density gradient-driven TEM one (pTEM). Some of the most important input parameters used in the simulations are reported in table 1, while the shapes of the three original scenarios with their flipped counterparts are shown in figure 1. For the realistic TCV and DIII-D simulations collisions, electromagnetic effects and carbon as the main impurity were retained. The pITG and pTEM simulations neglected impurities, but collisions and electromagnetic effects were kept. All simulations used a kinetic treatement for the electrons.
@@
Heat flux spectra are plotted in figure 3 for different regimes using two values of aspect ratio: conventional and spherical. We are not showing the spectra for the large aspect ratio cases because they are similar to conventional A. We show only the most important component, i.e. electron or ion channel, depending on the considered regime. From figures 3(a), 3(b) and 3(c), which correspond to ITG regimes, we can see that at conventional $A$ the reduction of heat flux is equally distributed among all the considered $k_{y}$ modes. Instead, when we move to a smaller $A$, the reduction is more pronounced for modes with $k_{y}\rho_{s}$ bigger than 0.2. If we now look at figure 3(d), which corresponds to the realistic TEM case, we can see that the detrimental effect of NT comes from small $k_{y}$ modes at tight A. These are the modes that are most strongly destabilized when one decreases the aspect ratio for TEM turbulence.